# A Taxonomy of Multi-Layered Runtime Guardrails for Designing Foundation Model-Based Agents: Swiss Cheese Model for AI Safety by Design

### v2

@@ Abstract
Foundation Model (FM)-based agents are revolutionizing application development across various domains. However, their rapidly growing capabilities and autonomy have raised significant concerns about AI safety. Designing effective guardrails for these agents is challenging due to their autonomous and non-deterministic behavior, and the involvement of multiple artifacts — such as goals, prompts, plans, tools, knowledge bases, and intermediate and final results. Addressing these unique challenges runtime requires multi-layered guardrails that operate effectively at various levels of the agent architecture, similar to the Swiss Cheese Model. In this paper, we present a taxonomy of multi-layered runtime guardrails to classify and compare their characteristics and design options, grounded on a systematic literature review and guided by the Swiss Cheese Model. This taxonomy is organized into external and internal quality attributes and design options categories. We also highlight the relationships between guardrails, the associated risks they mitigate, and the quality attributes they impact in agent architectures. Thus, the proposed taxonomy provides structured and concrete guidance for making architectural design decisions when implementing multi-layered guardrails while emphasizing the trade-offs inherent in these decisions.
@@ I Introduction
A foundation model (FM) is a large-scale machine learning model pre-trained on massive amounts of data using self-supervision at scale. These models are designed to be highly versatile and can adapt to a wide range of downstream tasks. The term ‘foundation’ reflects their role as the fundamental base upon which many specialized models are built [1]. FMs possess emergent capabilities, meaning their behaviors and susceptibilities arise implicitly from the training data rather than being explicitly programmed. This allows them to perform tasks that were not originally anticipated during their initial training [2]. An FM-based agent utilizes an FM as a core component, interacting with other AI or non-AI components to perform tasks autonomously [3]. In recent years, FM-based agents have experienced extensive growth [4].
The autonomy of FM-based agents along with their enhance decision-making and task execution capabilities introduce significant concerns about responsible AI and AI safety, e.g., generating harmful or offensive content, producing dangerous or unintended outcomes, exhibiting discriminatory behavior, compromising user privacy, spreading misinformation, facilitating cyberattacks, etc [1, 6, 8]. To address these challenges, effective runtime guardrails are required to ensure the agent’s behavior is responsible and safe. Multi-layered runtime guardrails structured similarly to the Swiss Cheese Model that operate at various levels of the agent architecture further mitigate associated risks [76, 1].
There have been some initial efforts on guardrails, such as input filtering [9, 1], output modification [10, 11], adaptive fail-safes that can prevent harmful outputs [12, 13], real-time monitoring and detection [14, 15, 16, 17], continuous output validation to ensure adherence to ethical standards [18, 19, 20] etc. However, the existing runtime guardrails mainly focus on inputs and outputs of foundation models, which do not capture the complexity of FM-based agents. FM-based agents are complex AI systems [21] that integrate foundation models and non-AI components such as data retrievers and external tools. Guardrails need to be placed across different components, taking into account the operational and environmental context, much like the layers in the Swiss cheese model [75]. However, most current work primarily addresses the functional correctness of guardrails, and quality attributes such as customizability and interpretability are often not considered.
Therefore, in this paper, we present a taxonomy for multilayer guardrails from a software architecture perspective. The taxonomy classifies and compares the characteristics and design options of guardrails. There are three main categories in the taxonomy: the external qualities for adopting multi-layered runtime guardrails, the internal attributes that should be addressed in the guardrails design, and the design options available. We developed this taxonomy based on the results of a systematic literature review (SLR), guided by the Swiss Cheese Model structure for AI safety. The main contributions of this paper are as follows:
  • The taxonomy offers a comprehensive comparative analysis of different design options and provides structured and concrete guidance for making architectural design decisions about multi-layered runtime guardrails to achieve AI-safety-by-design. The design options are classified into actions, targets, scopes, rules, autonomy, modalities, and underlying techniques.
  • This taxonomy helps researchers and practitioners understand the underlying rationales and benefits by summarizing the external qualities for multi-layered runtime guardrails. Hence, it also promotes informed decision-making when integrating guardrails into FM-based agents.
  • The taxonomy identify critical internal quality attributes to ensure that the multi-layered runtime guardrails not only perform their intended functions but also enhance the overall quality of the guardrails’ design.
The rest of the paper is organized as follows: section II discusses background and related works. section III introduces the methodology of this study. The taxonomy of guardrails is presented in section IV answering three key research questions. section V presents the implications of our proposed work and summarizes primary threats to this study. Finally, section VI concludes the paper and outlines the future work.
@@ II Background and Related Work
Foundation models (FMs) have significantly impacted the architecture of AI agents [1, 20]. Although guardrails for FM-based agents have been explored in various contexts, this paper is the first study assessing existing work to provide a comprehensive understanding. Below, we present key background and related works required to understand guardrails and the research gap.
II-A Background
The terms foundation models (FM) and large language models (LLMs) are often used interchangeably in the literature. However, they are not the same [22, 23]. Foundation models are large-scale machine learning models pre-trained on massive amounts of diverse datasets (such as text, images, videos) using self-supervised learning techniques [6]. They can perform a wide range of tasks across various application domains, including multimodal tasks. LLMs are a subset of FMs focused on language-related tasks and analyze data to perform natural language processing (NLP) tasks [22, 24]. Both FMs and LLMs are pre-trained on large datasets [25] and adapted for specific tasks using techniques like fine-tuning, in-context learning, distillation, and Retrieval-Augmented Generation (RAG) to improve response quality [24].
FM-based agents incorporate FMs as core components within larger architectures designed to utilize their capabilities [3]. These agents integrate FMs with additional components such as data retrievers (e.g., RAG), external tools and others. The complexity of FM-based agents lies in ensuring seamless interoperability among these components while ensuring responsible AI and AI safety [1].
@@ II-A Background
In 2021, Bommasani et al.[1] provided a comprehensive discussion on foundation models, illustrating the essential elements and relationships between them. The authors also discussed the practical applications of FM-based agents, highlighting their economic and operational advantages over traditional systems. Two years later, Zhou et al. covered recent research advancements, challenges, and opportunities for pre-trained foundation models in text, image, graph, and other data modalities in their survey paper[20]. Both works offer excellent insights into future research directions to address open problems and associated risks. Studies [1, 26] mention that foundation model-based agents often inherit responsible AI and AI safety issues and can cause negative consequences.
@@
Lu et al. developed a taxonomy of FM-based systems (agents) focusing on their pretraining, adaptation, architectural design, and responsible-AI-by-design [27]. The taxonomy aids software architects and developers in evaluating and integrating foundation models into complex systems. The authors also highlighted considerations for cost, accuracy, and responsible AI related quality attributes for the systems. The authors later proposed a reference architecture for designing responsible and safe foundation model based systems (agents) in [3]. The works [28, 29] explore the risks associated with deploying LLMs and evaluate current approaches to mitigate such risks through model alignment. Recently, Lu et al. discussed that responsible AI practices are highly recommended for FM-based systems (agents) [30, 2]. They also emphasized that FM-based agents must ensure they are developed and deployed in a responsible, safe, and legal way [32, 33, 34, 31].
II-B Related Work
There exist several frameworks and tools for designing guardrails at design time and runtime. Significant effort have been put into model alignment during design time, which focuses on aligning the FM’s outputs with human values and ethical standards. Pre-training and adaptation strategies play a significant role in mitigating risks in FM-based agents through guardrails, especially multilayer guardrails. Pre-training models on diverse and ethically sourced datasets helps align them more closely with ethical standards. Continuous adaptation through fine-tuning further reduces AI risks [1]. Ensuring ethical and responsible AI practices involves not only technical solutions but also adherence to regulatory and ethical guidelines within the systems. The adoption of comprehensive AI frameworks that incorporate ethical principles into the design and deployment of FMs is crucial for maintaining public trust and safety [31].
@@ II-B Related Work
Some initial efforts have been made toward runtime guardrails. Zhou et al. [20] present an excellent discussion of the literature that addresses the monitoring and controlling behavior of FM-based agents. Various tools and frameworks have also been developed to tackle these challenges. NeMo Guardrails[16] provides programmable guardrails to ensure that models operate within safe parameters by monitoring inputs and outputs in real-time, dynamically adjusting behavior to prevent harmful outcomes[27]. OpenAI’s Moderation API [35] monitors and filters harmful content generated by models to protect user interactions. The GuardAgent framework [36] utilizes an LLM agent to oversee and safeguard other agents. It demonstrates strong generalization and low operational overhead by dynamically generating guardrail code.
We also found that the continuous validation ensures that the outputs from FM-based agents adhere to predefined ethical standards and guidelines. Techniques such as auditing models [18] through multi-layered approaches [18, 37] systematically check for biases. They also ensure ethical compliance during the agent’s operation [3]. Dynamic access controls adjust access permissions in real time based on the context of data usage to protect sensitive information and ensure it is accessible under appropriate circumstances [38]. Adaptive fail-safes as characteristics of guardrails intervene automatically when an FM exhibits potentially harmful behavior within agents. These fail-safes are designed to modify or halt outputs that could lead to undesirable consequences [31]. However, the design of multi-layered runtime guardrails for FM-based agents has not been systematically explored. Thus, we develop a taxonomy to categorize existing work into a compact framework and compare and evaluate different design options for multi-layered runtime guardrails.
@@ III Methodology
This study focuses on two primary concepts: (i) foundation model-based agents and (ii) runtime guardrails. We adopted the Petticrew and Roberts approach [39] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [40], we develop the protocol for this study. The following subsections detail the steps and processes undertaken in this study.
@@
Population Studies and researches focus on runtime guardrails within foundation model-based agents.
Intervention Development, optimization, and evaluation of multilayer runtime guardrails in foundation model-based agents, focusing on key quality attributes and design strategies inspired by the Swiss Cheese Model.
Comparison Comparative analysis between different approaches to design multilayer runtime guardrails in FM-based agents.
Outcomes Qualities and design options for multilayer runtime guardrails for foundation model-based agents guided by Swiss cheese model.
@@ III-A Research Scope and Protocol Development
The high-level research approach for this study is shown in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [40, 41]. The protocol guided the entire study by defining relevant scientific databases and resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specifying criteria for the inclusion and exclusion of studies.
Figure 2: Study Selection Process for this SLR
@@ III-B Research Questions
When formulating our Research Questions (RQ), we wanted to ensure that they were broad enough to capture the diverse aspects of runtime guardrails while being specific enough to provide actionable insights. We captured these diverse aspects through the following three Research Questions (RQs).
RQ1: External qualities for runtime multilayer guardrails in FM-based agents?
Our first research question investigates the fundamental external qualities for designing multilayer runtime guardrails in FM-based agents. It explores the responsible AI and AI safety concerns that require the adoption of runtime guardrails to prevent harmful content and unintended behaviors.
RQ2: What are the key internal quality attributes for designing multilayer runtime guardrails?
Our second research question identifies quality attributes that should be considered when designing effective guardrails in FM-based agents.
RQ3: What are the design options for runtime guardrails?
This question examines the design options for runtime guardrails in FM-based agents from different perspectives, such as action and scope, guided by the Swiss Cheese Model.
@@ III-C Search String Formulation
[…]
$\textbf{NOT}(UC_{1}\|UC_{2}\|...\|UC_{n})]$
where $C_{11...1n}$, and $C_{21...2n}$ $\varepsilon$ Co1 and Co2 of Table II, respectively; and $UC_{1}$ $UC_{n}$ refers the Exclude Context defined earlier in PICOC table (Table I).
III-D Selection of Papers: Inclusion and Exclusion Criterion
@@ III-E Study Search and Filtering Process
Our filtration process is further detailed in Figure 2. Initially we ran the formatted query on four major databases that returned 1,733 research papers. We then applied filtering and classified the studies found according to the guidelines presented in [41, 40]. In our initial filtration process, we removed 108 papers due to being duplicated articles, editorial or key notes. After reading the title, abstract, conclusion and skimming through the introduction, methodology and results, we applied our exclusion criterion defined in Table IV, and 1524 further papers were removed. During the third step of filtration, we applied inclusion criteria and removed 80 papers as these studies did not meet ICs shown in Table IV. In parallel, we did a manual search and found 189 papers that meet our key concepts defined in Table II but not contain any unwanted content (UC). After applying ICs and ECs, 15 out of 189 papers were selected. Finally, we did a cross-check and ended up with 32 papers.
@@ III-F Data Extraction and Quality Assessment
We then evaluated each study based on the following five Quality Assessment Criteria (QAC) on a scale from 1 (Very Poor) to 5 (Excellent). If a study’s average score was less than 2, it was excluded from further analysis. Otherwise, we used the qualitative information to decide this. The QAC for this study are as follows:
@@
IV The Proposed Taxonomy
In this section, we present our taxonomy of multilayer runtime guardrails for FM-based agents. As illustrated in Figure 3, this taxonomy is organized into three main categories: external quality attributes, internal quality attributes, and design options, which we use in the following sub-sections to answer our key research questions guided by the Swiss Cheese Model for AI Safety by Design.
IV-A External Qualities for Multi-layered Runtime Guardrails (RQ1)
Our first research question investigates the fundamental reasons for integrating guardrails into FM-based agents. We found that guardrails are primarily implemented to mitigate bias, harmful content, and unintended behaviors. Additionally, guardrails better maintain ethical and safety standards in areas such as accuracy, privacy, security, safety, fairness, cost, performance and IP protection, as discussed below.
Figure 3: Taxonomy of Multi-Layered Runtime Guardrails for FM-Based Agents

### v3

# Designing Multi-layered Runtime Guardrails for Foundation Model Based Agents: Swiss Cheese Model for AI Safety by Design

@@ Abstract
Foundation Model (FM)-based agents are revolutionizing application development across various domains. However, their rapidly growing capabilities and autonomy have raised significant concerns about AI safety. Researchers are exploring better ways to design guardrails to ensure that the runtime behavior of FM-based agents remains within specific boundaries. Nevertheless, designing effective runtime guardrails is challenging due to the agents’ autonomous and non-deterministic behavior. The involvement of multiple pipeline stages and agent artifacts, such as goals, plans, tools, at runtime further complicates these issues. Addressing these challenges at runtime requires multi-layered guardrails that operate effectively at various levels of the agent architecture. Thus, in this paper, we present a comprehensive taxonomy of runtime guardrails for FM-based agents to identify the key quality attributes for guardrails and design dimensions based on the results of a systematic literature review. Inspired by the Swiss Cheese Model, we also propose a reference architecture for designing multi-layered runtime guardrails for FM-based agents, which includes three dimensions: quality attributes, pipelines, and artifacts. The proposed taxonomy and reference architecture provide concrete and robust guidance for researchers and practitioners to build AI-safety-by-design from a software architecture perspective.
@@ I Introduction
A Foundation Model (FM) is a large-scale machine learning model pre-trained on massive amounts of data using self-supervision at scale. These models are highly versatile and can adapt to a wide range of downstream tasks [1]. The term ‘foundation’ reflects their role as the fundamental base upon which many specialized models/systems are built. However, it is important to recognize that FM-based systems exhibit inherent limitations, particularly when handling complex tasks. Users are often required to provide detailed instructions, which can lead to inefficiencies and is prone to error.
An FM-based agent is an autonomous system that is capable of perceiving context, reasoning, planning, and executing workflows by interacting with FMs, external tools, knowledge bases, and other agents to achieve human goals [3]. There has been extensive interest in FM-based agent development recently due to their huge potential to enhance productivity across various domains. However, their autonomous and non-deterministic behavior introduce substantial concerns regarding AI safety [2, 78], such as generating harmful or offensive content, producing dangerous or unintended outcomes, spreading disinformation and misinformation, etc [77].
To address these challenges, effective runtime guardrails are key to ensure that agents behave in a safe and responsible manner [2]. In this context, guardrails are mechanisms integrated into the agent’s architecture to safeguard its behavior during runtime, preventing undesirable or unsafe behaviors [78]. There have been some initial efforts on runtime guardrails such as input filtering [9, 1], output modification [10, 11], adaptive fail-safes [12, 13], real-time monitoring and detection [14, 15, 16, 17], and continuous output validation [18, 19, 20].
However, the existing guardrail approaches primarily address functional correctness, often overlooking quality attributes of FM-based agents, such as customizability and interpretability. Most importantly, these approaches mainly focus on individual single-layered guardrails that are narrowly applied to specific agent artifacts, such as prompts or FM outputs, which are insufficient to manage the inherent autonomy and non-deterministic nature of FM-agents. If any single guardrail fails, the associated risks may bypass it, potentially impacting the final results of the FM-based agent.
Therefore, in this paper, we first present a comprehensive taxonomy to categorize runtime guardrails from a software architecture perspective, based on the results of a systematic literature review. The taxonomy comprises two primary categories: quality attributes and design options. Inspired by Swiss Cheese Model [76], we also propose novel reference architecture for designing multi-layered guardrails of FM-based agents which include three dimensions: quality attributes, pipelines, and artifacts. Each guardrail layer can be designed to protect specific quality attributes (such as privacy and security), specific pipeline stages (such as prompts, intermediate results and final results), as well as agent artifacts (such as goals, plans, and tools). While each layer may have its own weaknesses (i.e. holes in the Swiss Cheese Model), the combined layers create a a robust defense against failures. This reference architecture provides concrete guidance for researchers and practitioners, enabling AI-safety-by-design from a software architecture perspective.
The rest of the paper is organized as follows. section II discuss the related works and background study required to understand the proposed works. The research methods employed in this study are described in section III, including a brief discussion of the research protocol used for systematic literature review. The proposed taxonomy of guardrails is presented in section IV,developed based on the results of a systematic literature review. The taxonomy is organized into guardrails quality attributes and design options from different perspectives. section V proposes the reference architecture for multi-layered runtime guardrails for FM-based agents. section VI identifies and summarizes the primary threats that could impact the validity of this study. Finally, section VII concludes the paper and outlines directions for future work.
@@ II Background and Related Work
FMs have significantly advanced current agent development and emphasize the need to safeguard their behavior [3, 20]. In this context, guardrails for FM-based agents have been explored; however, there is a lack of comprehensive studies that provide a thorough understanding of guardrails for FM-based agents. This paper aims to fill this gap. In the following sections, we present key background information and related work.
II-A Recent State-of-the-Art Works on Foundation Models and FM-Based Agents
@@
In 2021, Bommasani et al. [1] provided a comprehensive discussion on FMs, illustrating key elements, relationships, opportunities, and associated risks. While their focus was on FMs in general, they highlighted the potential for these models to serve as the foundation for more complex systems, including FM-based agents. Zhou et al. [20] reviewed research advancements, challenges, and opportunities for pre-trained models in text, image, graph, and data modalities. They also discussed the integration of FMs into systems such as agents. Both works offer excellent insights into future research directions to address open problems and associated risks in FM-based agents.
@@
Recently, Lu et al. developed a taxonomy of FM-based systems focusing on their pre-training, adaptation, architectural design, and responsible-AI-by-design [27]. The taxonomy aids software architects and developers in evaluating and integrating FMs into complex agent systems. The authors then highlighted considerations for responsible AI and safety attributes. Several other works [30, 2, 31, 36, 8], also emphasize the importance of responsible AI and safety practices for FM-based agents. In [28, 29], the authors explored the risks associated with deploying LLM-based agents and evaluated current approaches for mitigating these risks through model alignment, respectively. In 2024, a reference architecture for designing responsible and safe FM-based agents is proposed in [3]. The authors demonstrated that the unique characteristics of FM-based agents—such as their autonomous operation, non-deterministic behavior, and continuous evolution—pose significant challenges in ensuring responsible AI and AI safety.
II-B Existing Guardrails Approaches and Tools for FM-Based Agents
There exist several frameworks and tools for designing guardrails [55, 67, 9, 53, 62]. These works explored model alignment during design time to ensure that the FM’s outputs align with defined goals. Pre-training and adaptation strategies play a significant role in mitigating risks in FM-based agents. Our focus, however, is on runtime guardrails that monitor and control the agent’s behavior during operation. These guardrails are essential for addressing emergent issues that arise during agent interactions within dynamic environments [1, 31].
@@ II-B Existing Guardrails Approaches and Tools for FM-Based Agents
Some initial efforts have been made toward runtime guardrails. NeMo Guardrails [16] provides programmable guardrails to ensure that agents operate within safe parameters by monitoring inputs and outputs. OpenAI’s Moderation API [35] monitors and filters harmful content generated by agents to protect user interactions. The GuardAgent framework [36] utilizes an agent to oversee and safeguard other agents. It demonstrates strong generalization and low operational overhead by dynamically generating guardrail code. We found that continuous validation ensures outputs from FM-based agents adhere to predefined ethical standards and guidelines. Techniques such as auditing agents through multi-layered approaches [18, 37] are used to check for biases and ensure ethical compliance.
Recently, Bengio et. al. [31] demonstrate that adaptive fail-safes characteristics of guardrails intervene automatically when an FM-based agent exhibits potentially harmful behavior. These fail-safes are designed to modify or halt outputs that could lead to undesirable consequences. Similarly, dynamic access controls adjust access permissions in real time based on the context of data usage to protect sensitive information and ensure it is accessible under appropriate circumstances [38]. Due to the dynamic and adaptive nature of FM-based agents, designing effective runtime guardrails poses several additional challenges [2, 8] e.g., scalability of guardrail mechanisms, the need for real-time monitoring, and the complexity of interpreting agent behaviors in diverse contexts. The authors in [77] propose a framework for evaluating AI systems, which is applicable to FM-based agents. It includes harmonized terminology, a taxonomy of key evaluation elements, and a mapping of the AI lifecycle to stakeholders for ethical and accountable deployment. Despite these efforts, no framework currently provides comprehensive guidance on designing multi-layered runtime guardrails for FM-based agents, which we explore in this paper based on SLR.
@@ III Methodology
This study focuses on two primary concepts: (i) foundation model-based agents and (ii) runtime guardrails. We adopted the Petticrew and Roberts approach [39] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [40], we develop the protocol for this study.
@@
Population Studies and researches focus on multi-layered runtime guardrails within foundation model-based agents.
Intervention Development, optimization, and evaluation of multilayer runtime guardrails in foundation model-based agents, focusing on key quality attributes and design strategies similar to the Swiss Cheese Model structure.
Comparison Comparative analysis of approaches to design multi-layered runtime guardrails in FM-based agents.
Outcomes Taxonomy of multi-layered runtime guardrails for foundation model-based agents.
@@ III-A Research Scope and Protocol Development
The high-level research approach for this study is shown in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [40, 41]. The protocol guided the entire study by defining relevant scientific databases and resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specify criteria for study inclusion and exclusion.
@@ III-B Research Questions
When formulating our Research Questions (RQ), we wanted to ensure that they were broad enough to capture the diverse aspects of multi-layered runtime guardrails while being specific enough to provide actionable insights. We captured these aspects through the following three RQs:
RQ1: What are essential qualities for designing runtime guardrails in FM-based agents?
Our first research question studies the key qualities for designing multi-layered runtime guardrails in FM-based agents. section IV-A elaborate on how this research question is addressed.
RQ2: What are the design options for runtime guardrails in FM-based agents?
Our second research question investigates guardrails design options in FM-based agents from different perspectives, including action, target, scope, rule, autonomy, modalities, and underlying techniques. section IV-B outlines our approach to addressing this research question.
RQ3: How can we design runtime guardrails to address the unique challenges of FM-based agents?
Our third research question explores how to address the challenges arising from the autonomous and deterministic nature of FM-based agents. Specially, we examine how to adapt the Swiss Cheese Model to safeguard the behaviors of FM-based agents by implementing multi-layered guardrails across various agent artifacts. section V presents the proposed architecture and discusses our strategies for addressing this research question.
@@ III-C Search String Formulation
[…]
$\textbf{NOT}(UC_{1}\|UC_{2}\|...\|UC_{n})]$
where $C_{11,12,...,1n}$, and $C_{21,22,...,2n}$ $\varepsilon$ Co1 and Co2 of Table II, respectively; and $UC_{1}$, $UC_{2}$, . . ., $UC_{n}$ refers the Exclude Context defined earlier in PICOC (Table I).
III-D Selection of Papers: Inclusion and Exclusion Criterion
@@ III-E Study Search and Filtering Process
Our filtration process is further detailed in Figure 2. Initially we ran the formatted query on four major databases that returned 1,733 research papers. We then applied filtering and classified the studies found according to the guidelines presented in [41, 40]. In our initial filtration process, we removed 108 papers due to being duplicated articles, editorial or key notes. After reading the title, abstract, conclusion and skimming through the introduction, methodology and results, we applied our exclusion criterion defined in Table IV, and 1524 further papers were removed. During the third step of filtration, we applied inclusion criteria and removed 80 papers as these studies did not meet ICs shown in Table IV. In parallel, we did a manual search and found 189 papers that meet our key concepts defined in Table II but not contain any unwanted content (UC). After applying ICs and ECs, 15 out of 189 papers were selected. Finally, we did a cross-check and ended up with 32 papers (shown in Appendix A).
Figure 2: Study Selection Process for this SLR
@@ III-F Data Extraction and Quality Assessment
We then evaluated each study based on the following five Quality Assessment Criteria (QAC) on a scale from 1 (Very Poor) to 5 (Excellent). If a study’s average score was less than 2, it was excluded from further analysis. Otherwise, we used the qualitative information to decide this. The QAC used for this study are  [^1] QAC score for each selected study is presented in Appendix B:
@@
Figure 3: Taxonomy of multi-layered runtime guardrails for FM-based agents.
IV Taxonomy of Guardrails for FM-based Agents
Figure 3 presents the proposed taxonomy of runtime guardrails for FM-based agents, developed based on the results of a systematic literature review. The taxonomy is organized into external and internal quality attributes, and design options from different perspectives.
IV-A Quality Attributes of Guardrails
We examine the key quality attributes that should be considered when designing runtime guardrails. Below, we discuss these attributes in detail.